# ADOPT: Modified Adam Can Converge with Any $\beta_{2}$ with the Optimal Rate

### v1

@@ Appendix G Lemmas
$\mathbb{E}\left[\sum_{i=1}^{D}\left({\bm{v}}_{t}\right)_{i}\right]$ $=\mathbb{E}\left[\left(1-\beta_{2}\right)\sum_{i=1}^{D}\sum_{k=1}^{t}\beta_{2}^{t-k}\left({\bm{g}}_{k-1}\right)_{i}^{2}\right]$ (121)
$\leq\left(1-\beta_{2}\right)G^{2}\sum_{k=1}^{t}\beta_{2}^{t-k}$ (122)
$=\left(1-\beta_{2}^{t}\right)G^{2}.$ (123)
@@
$\mathbb{E}\left[\left\|{\bm{m}}_{t}\right\|^{2}\right]\leq\frac{G^{2}}{\epsilon^{2}}.$ (124)
@@
$=\mathbb{E}\left[\left\|\beta_{1}{\bm{m}}_{t-1}+\left(1-\beta_{1}\right)\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}\right\|^{2}\right]$ (125)
$=\mathbb{E}\left[\beta_{1}^{2}\left\|{\bm{m}}_{t-1}\right\|^{2}+\left(1-\beta_{1}\right)^{2}\left\|\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}\right\|^{2}+2\beta_{1}\left(1-\beta_{1}\right){\bm{m}}_{t-1}^{\top}\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}\right]$ (126)
$\leq\mathbb{E}\left[\beta_{1}\left\|{\bm{m}}_{t-1}\right\|^{2}+\left(1-\beta_{1}\right)\left\|\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}\right\|^{2}\right]$ (127)
$\leq\mathbb{E}\left[\beta_{1}\left\|{\bm{m}}_{t-1}\right\|^{2}+\frac{1-\beta_{1}}{\epsilon^{2}}\left\|{\bm{g}}_{t}\right\|^{2}\right]$ (128)
$\leq\mathbb{E}\left[\frac{1-\beta_{1}}{\epsilon^{2}}\sum_{k=1}^{t}\beta_{1}^{t-k}\left\|{\bm{g}}_{k}\right\|^{2}\right]$ (129)
$\leq\frac{\left(1-\beta_{1}\right)G^{2}}{\epsilon^{2}}\sum_{k=1}^{t}\beta_{1}^{t-k}$ (130)
$\leq\frac{\left(1-\beta_{1}^{t}\right)G^{2}}{\epsilon^{2}}$ (131)
$\leq\frac{G^{2}}{\epsilon^{2}}.$ (132)
@@
By setting $\lambda=\left(1-\beta_{1}\right)/\beta_{1},{\bm{x}}=\beta_{1}{\bm{m}}_{t-1},{\bm{y}}=\left(1-\beta_{1}\right){\bm{g}}_{t}/\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}$, we obtain
$2\beta_{1}\left(1-\beta_{1}\right){\bm{m}}_{t-1}^{\top}\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}\leq\beta_{1}\left(1-\beta_{1}\right)\left(\left\|{\bm{m}}_{t-1}\right\|^{2}+\left\|\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}\right\|^{2}\right)$ (134)
@@
$\mathbb{E}\left[\left\|{\bm{m}}_{t}\right\|\right]\leq\frac{G}{\epsilon}$ (135)
@@
$\mathbb{E}\left[\left\|{\bm{m}}_{t}\right\|\right]$ $=\mathbb{E}\left[\left\|\left(1-\beta_{1}\right)\sum_{k=1}^{t}\beta_{1}^{t-k}\frac{{\bm{g}}_{k}}{\max\left\{\sqrt{{\bm{v}}_{k-1}},\epsilon\right\}}\right\|\right]$ (136)
$\leq\left(1-\beta_{1}\right)\sum_{k=1}^{t}\beta_{1}^{t-k}\mathbb{E}\left[\left\|\frac{{\bm{g}}_{k}}{\max\left\{\sqrt{{\bm{v}}_{k-1}},\epsilon\right\}}\right\|\right]$ (137)
$\leq\left(1-\beta_{1}\right)\sum_{k=1}^{t}\frac{\beta_{1}^{t-k}}{\epsilon}\mathbb{E}\left[\left\|{\bm{g}}_{k}\right\|\right]$ (138)
$\leq\frac{1-\beta_{1}}{\epsilon}\sum_{k=1}^{t}\beta_{1}^{t-k}\mathbb{E}\left[\left\|{\bm{g}}_{k}\right\|^{2}\right]^{1/2}$ (139)
$\leq\frac{\left(1-\beta_{1}\right)G}{\epsilon}\sum_{k=1}^{t}\beta_{1}^{t-k}$ (140)
$=\frac{\left(1-\beta_{1}^{t}\right)G}{\epsilon}$ (141)
$\leq\frac{G}{\epsilon}.$ (142)
@@
$\leq\frac{\alpha_{t-1}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}^{2}\left(1-\beta_{1}^{t-1}\right)G^{2}}{\epsilon^{2}\left(1-\beta_{1}\right)^{2}}+\frac{\alpha_{t}\alpha_{t-1}\beta_{1}\sqrt{1-\beta_{1}^{t-1}}G^{2}}{\epsilon^{2}\left(1-\beta_{1}\right)}.$ (143)
@@
$=\left\|-\frac{\alpha_{t-1}\beta_{1}}{1-\beta_{1}}{\bm{m}}_{t-1}\right\|\left\|\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}}{1-\beta_{1}}{\bm{m}}_{t-1}-\alpha_{t}\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}\right\|$ (144)
$\leq\frac{\alpha_{t-1}\beta_{1}}{1-\beta_{1}}\left\|{\bm{m}}_{t-1}\right\|\left(\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}}{1-\beta_{1}}\left\|{\bm{m}}_{t-1}\right\|+\alpha_{t}\left\|\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}\right\|\right)$ (145)
$\leq\frac{\alpha_{t-1}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}^{2}}{\left(1-\beta_{1}\right)^{2}}\left\|{\bm{m}}_{t-1}\right\|^{2}+\frac{\alpha_{t}\alpha_{t-1}\beta_{1}}{1-\beta_{1}}\left\|{\bm{m}}_{t-1}\right\|\left\|\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}\right\|.$ (146)
@@
$\leq\frac{\alpha_{t-1}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}^{2}}{\left(1-\beta_{1}\right)^{2}}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|^{2}\right]+\frac{\alpha_{t}\alpha_{t-1}\beta_{1}}{1-\beta_{1}}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|\left\|\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}\right\|\right]$ (147)
$\leq\frac{\alpha_{t-1}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}^{2}}{\left(1-\beta_{1}\right)^{2}}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|^{2}\right]+\frac{\alpha_{t}\alpha_{t-1}\beta_{1}}{\left(1-\beta_{1}\right)\epsilon}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|\left\|{\bm{g}}_{t}\right\|\right]$ (148)
$\leq\frac{\alpha_{t-1}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}^{2}\left(1-\beta_{1}^{t-1}\right)G^{2}}{\epsilon^{2}\left(1-\beta_{1}\right)^{2}}+\frac{\alpha_{t}\alpha_{t-1}\beta_{1}\sqrt{1-\beta_{1}^{t-1}}G^{2}}{\left(1-\beta_{1}\right)\epsilon^{2}}.$ (149)
@@
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)^{2}\beta_{1}^{2}\left(1-\beta_{1}^{t-1}\right)G^{2}}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}}+\frac{\alpha_{t}^{2}G^{2}}{\epsilon^{2}}+\frac{\alpha_{t}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}\sqrt{1-\beta_{1}^{t-1}}G^{2}}{\left(1-\beta_{1}\right)\epsilon^{2}}.$ (150)
@@
$=\left\|\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}}{1-\beta_{1}}{\bm{m}}_{t-1}-\alpha_{t}\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}\right\|^{2}$ (151)
$=\frac{\left(\alpha_{t-1}-\alpha_{t}\right)^{2}\beta_{1}^{2}}{\left(1-\beta_{1}\right)^{2}}\left\|{\bm{m}}_{t-1}\right\|^{2}+\alpha_{t}^{2}\left\|\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}\right\|^{2}$
$\quad-\frac{\alpha_{t}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}}{1-\beta_{1}}{\bm{m}}_{t-1}^{\top}\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}$ (152)
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)^{2}\beta_{1}^{2}}{\left(1-\beta_{1}\right)^{2}}\left\|{\bm{m}}_{t-1}\right\|^{2}+\frac{\alpha_{t}^{2}}{\epsilon^{2}}\left\|{\bm{g}}_{t}\right\|^{2}$
$\quad+\frac{\alpha_{t}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}}{1-\beta_{1}}\left\|{\bm{m}}_{t-1}\right\|\left\|\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}\right\|$ (153)
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)^{2}\beta_{1}^{2}}{\left(1-\beta_{1}\right)^{2}}\left\|{\bm{m}}_{t-1}\right\|^{2}+\frac{\alpha_{t}^{2}}{\epsilon^{2}}\left\|{\bm{g}}_{t}\right\|^{2}+\frac{\alpha_{t}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}}{\left(1-\beta_{1}\right)\epsilon}\left\|{\bm{m}}_{t-1}\right\|\left\|{\bm{g}}_{t}\right\|.$ (154)
@@
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)^{2}\beta_{1}^{2}}{\left(1-\beta_{1}\right)^{2}}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|^{2}\right]+\frac{\alpha_{t}^{2}}{\epsilon^{2}}\mathbb{E}\left[\left\|{\bm{g}}_{t}\right\|^{2}\right]+\frac{\alpha_{t}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}}{\left(1-\beta_{1}\right)\epsilon}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|\left\|{\bm{g}}_{t}\right\|\right]$ (155)
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)^{2}\beta_{1}^{2}\left(1-\beta_{1}^{t-1}\right)G^{2}}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}}+\frac{\alpha_{t}^{2}G^{2}}{\epsilon^{2}}+\frac{\alpha_{t}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}}{\left(1-\beta_{1}\right)\epsilon}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|\left\|{\bm{g}}_{t}\right\|\right]$ (156)
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)^{2}\beta_{1}^{2}\left(1-\beta_{1}^{t-1}\right)G^{2}}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}}+\frac{\alpha_{t}^{2}G^{2}}{\epsilon^{2}}+\frac{\alpha_{t}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}\sqrt{1-\beta_{1}^{t-1}}G^{2}}{\left(1-\beta_{1}\right)\epsilon^{2}}.$ (157)

### v3

@@ Appendix G Lemmas
$\mathbb{E}\left[\sum_{i=1}^{D}\left({\bm{v}}_{t}\right)_{i}\right]$ $=\mathbb{E}\left[\beta_{2}^{t}\sum_{i=1}^{D}\left({\bm{g}}_{0}\right)_{i}^{2}+\left(1-\beta_{2}\right)\sum_{i=1}^{D}\sum_{k=1}^{t}\beta_{2}^{t-k}\left({\bm{g}}_{k-1}\right)_{i}^{2}\right]$ (150)
$\leq\beta_{2}^{t}G^{2}+\left(1-\beta_{2}\right)G^{2}\sum_{k=1}^{t}\beta_{2}^{t-k}$ (151)
$=\beta_{2}^{t}G^{2}+\left(1-\beta_{2}^{t}\right)G^{2}$ (152)
$=G^{2}$ (153)
Lemma G.12.
For the ADOPT algorithm, the following holds for $t\geq 0$:
$\mathbb{E}\left[\left\|\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)\right\|^{2}\right]\leq\frac{2G^{2}}{\epsilon^{2}}.$ (154)
Proof.
$\mathbb{E}\left[\left\|\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)\right\|^{2}\right]$ (155)
$=\mathbb{E}\left[\sum_{i=1}^{D}\mathrm{Clip}\left(\frac{\left({\bm{g}}_{t}\right)_{i}}{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}},c_{t}\right)^{2}\right]$ (156)
$=\mathbb{E}\left[\sum_{i=1}^{D}\delta_{t,i}c_{t}^{2}+\left(1-\delta_{t,i}\right)\frac{\left({\bm{g}}_{t}\right)_{i}^{2}}{\max\left\{\left({\bm{v}}_{t-1}\right)_{i},\epsilon^{2}\right\}}\right]$ (157)
$\leq\sum_{i=1}^{D}c_{t}^{2}\mathbb{E}\left[\delta_{t,i}\right]+\mathbb{E}\left[\frac{\left({\bm{g}}_{t}\right)_{i}^{2}}{\max\left\{\left({\bm{v}}_{t-1}\right)_{i},\epsilon^{2}\right\}}\right]$ (158)
$=\sum_{i=1}^{D}c_{t}^{2}\mathrm{Pr}\left[\left|\frac{\left({\bm{g}}_{t}\right)_{i}}{\max\left\{\sqrt{\left({\bm{v}}_{t-1}\right)_{i}},\epsilon\right\}}\right|>c_{t}\right]+\mathbb{E}\left[\frac{\left({\bm{g}}_{t}\right)_{i}^{2}}{\max\left\{\left({\bm{v}}_{t-1}\right)_{i},\epsilon^{2}\right\}}\right]$ (159)
$\leq 2\cdot\mathbb{E}\left[\sum_{i=1}^{D}\frac{\left({\bm{g}}_{t}\right)_{i}^{2}}{\max\left\{\left({\bm{v}}_{t-1}\right)_{i},\epsilon^{2}\right\}}\right]$ (160)
$\leq\frac{2}{\epsilon^{2}}\cdot\mathbb{E}\left[\left\|{\bm{g}}_{t}\right\|^{2}\right]$ (161)
$\leq\frac{2G^{2}}{\epsilon^{2}},$ (162)
where the second inequality is due to the Chebyshev’s inequality. ∎
@@
$\mathbb{E}\left[\left\|{\bm{m}}_{t}\right\|^{2}\right]\leq\frac{2G^{2}}{\epsilon^{2}}.$ (163)
@@
$=\mathbb{E}\left[\left\|\beta_{1}{\bm{m}}_{t-1}+\left(1-\beta_{1}\right)\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)\right\|^{2}\right]$ (164)
$=\mathbb{E}\left[\beta_{1}^{2}\left\|{\bm{m}}_{t-1}\right\|^{2}+\left(1-\beta_{1}\right)^{2}\left\|\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)\right\|^{2}\right]$
$\quad+\mathbb{E}\left[2\beta_{1}\left(1-\beta_{1}\right){\bm{m}}_{t-1}^{\top}\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)\right]$ (165)
$\leq\mathbb{E}\left[\beta_{1}\left\|{\bm{m}}_{t-1}\right\|^{2}+\left(1-\beta_{1}\right)\left\|\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)\right\|^{2}\right]$ (166)
$\leq\mathbb{E}\left[\beta_{1}\left\|{\bm{m}}_{t-1}\right\|^{2}+\frac{2\left(1-\beta_{1}\right)G^{2}}{\epsilon^{2}}\right]$ (167)
$\leq\frac{2\left(1-\beta_{1}\right)G^{2}}{\epsilon^{2}}\sum_{k=1}^{t}\beta_{1}^{t-k}$ (168)
$\leq\frac{2\left(1-\beta_{1}^{t}\right)G^{2}}{\epsilon^{2}}$ (169)
$\leq\frac{2G^{2}}{\epsilon^{2}}.$ (170)
@@
By setting $\lambda=\left(1-\beta_{1}\right)/\beta_{1},{\bm{x}}=\beta_{1}{\bm{m}}_{t-1},{\bm{y}}=\left(1-\beta_{1}\right)\mathrm{Clip}\left({\bm{g}}_{t}/\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\},c_{t}\right)$, we obtain
$2\beta_{1}\left(1-\beta_{1}\right){\bm{m}}_{t-1}^{\top}\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)$ (172)
$\leq\beta_{1}\left(1-\beta_{1}\right)\left(\left\|{\bm{m}}_{t-1}\right\|^{2}+\left\|\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)\right\|^{2}\right)$ (173)
@@
$\mathbb{E}\left[\left\|{\bm{m}}_{t}\right\|\right]\leq\frac{\sqrt{2}G}{\epsilon}$ (174)
@@
$\mathbb{E}\left[\left\|{\bm{m}}_{t}\right\|\right]$ $=\mathbb{E}\left[\left\|\left(1-\beta_{1}\right)\sum_{k=1}^{t}\beta_{1}^{t-k}\mathrm{Clip}\left(\frac{{\bm{g}}_{k}}{\max\left\{\sqrt{{\bm{v}}_{k-1}},\epsilon\right\}},c_{t}\right)\right\|\right]$ (175)
$\leq\left(1-\beta_{1}\right)\sum_{k=1}^{t}\beta_{1}^{t-k}\mathbb{E}\left[\left\|\mathrm{Clip}\left(\frac{{\bm{g}}_{k}}{\max\left\{\sqrt{{\bm{v}}_{k-1}},\epsilon\right\}},c_{t}\right)\right\|\right]$ (176)
$\leq\left(1-\beta_{1}\right)\sum_{k=1}^{t}\beta_{1}^{t-k}\sqrt{\mathbb{E}\left[\left\|\mathrm{Clip}\left(\frac{{\bm{g}}_{k}}{\max\left\{\sqrt{{\bm{v}}_{k-1}},\epsilon\right\}},c_{t}\right)\right\|^{2}\right]}$ (177)
$\leq\frac{\sqrt{2}\left(1-\beta_{1}\right)G}{\epsilon}\sum_{k=1}^{t}\beta_{1}^{t-k}$ (178)
$=\frac{\sqrt{2}\left(1-\beta_{1}^{t}\right)G}{\epsilon}$ (179)
$\leq\frac{\sqrt{2}G}{\epsilon}.$ (180)
@@
$\leq\frac{2\alpha_{t-1}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}^{2}\left(1-\beta_{1}^{t-1}\right)G^{2}}{\epsilon^{2}\left(1-\beta_{1}\right)^{2}}+\frac{2\alpha_{t}\alpha_{t-1}\beta_{1}\sqrt{1-\beta_{1}^{t-1}}G^{2}}{\epsilon^{2}\left(1-\beta_{1}\right)}.$ (181)
@@
$=\left\|-\frac{\alpha_{t-1}\beta_{1}}{1-\beta_{1}}{\bm{m}}_{t-1}\right\|\left\|\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}}{1-\beta_{1}}{\bm{m}}_{t-1}-\alpha_{t}\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)\right\|$ (182)
$\leq\frac{\alpha_{t-1}\beta_{1}}{1-\beta_{1}}\left\|{\bm{m}}_{t-1}\right\|\left(\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}}{1-\beta_{1}}\left\|{\bm{m}}_{t-1}\right\|+\alpha_{t}\left\|\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)\right\|\right)$ (183)
$\leq\frac{\alpha_{t-1}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}^{2}}{\left(1-\beta_{1}\right)^{2}}\left\|{\bm{m}}_{t-1}\right\|^{2}+\frac{\alpha_{t}\alpha_{t-1}\beta_{1}}{1-\beta_{1}}\left\|{\bm{m}}_{t-1}\right\|\left\|\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)\right\|.$ (184)
@@
$\leq\frac{\alpha_{t-1}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}^{2}}{\left(1-\beta_{1}\right)^{2}}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|^{2}\right]+\frac{\alpha_{t}\alpha_{t-1}\beta_{1}}{1-\beta_{1}}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|\left\|\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)\right\|\right]$ (185)
$\leq\frac{\alpha_{t-1}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}^{2}}{\left(1-\beta_{1}\right)^{2}}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|^{2}\right]+\frac{\sqrt{2}\alpha_{t}\alpha_{t-1}\beta_{1}G}{\left(1-\beta_{1}\right)\epsilon}\mathbb{E}\left[\left\|{\bm{m}}_{t-1}\right\|\right]$ (186)
$\leq\frac{2\alpha_{t-1}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}^{2}\left(1-\beta_{1}^{t-1}\right)G^{2}}{\epsilon^{2}\left(1-\beta_{1}\right)^{2}}+\frac{2\alpha_{t}\alpha_{t-1}\beta_{1}\sqrt{1-\beta_{1}^{t-1}}G^{2}}{\left(1-\beta_{1}\right)\epsilon^{2}}.$ (187)
@@
$\leq\frac{2\left(\alpha_{t-1}-\alpha_{t}\right)^{2}\beta_{1}^{2}\left(1-\beta_{1}^{t-1}\right)G^{2}}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}}+\frac{2\alpha_{t}^{2}G^{2}}{\epsilon^{2}}+\frac{2\alpha_{t}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}\sqrt{1-\beta_{1}^{t-1}}G^{2}}{\left(1-\beta_{1}\right)\epsilon^{2}}.$ (188)
@@
$=\left\|\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}}{1-\beta_{1}}{\bm{m}}_{t-1}-\alpha_{t}\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)\right\|^{2}$ (189)
$=\frac{\left(\alpha_{t-1}-\alpha_{t}\right)^{2}\beta_{1}^{2}}{\left(1-\beta_{1}\right)^{2}}\left\|{\bm{m}}_{t-1}\right\|^{2}+\alpha_{t}^{2}\left\|\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)\right\|^{2}$
$\quad-\frac{\alpha_{t}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}}{1-\beta_{1}}{\bm{m}}_{t-1}^{\top}\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)$ (190)
$\leq\frac{\left(\alpha_{t-1}-\alpha_{t}\right)^{2}\beta_{1}^{2}}{\left(1-\beta_{1}\right)^{2}}\left\|{\bm{m}}_{t-1}\right\|^{2}+\alpha_{t}^{2}\left\|\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)\right\|^{2}$
$\quad+\frac{\alpha_{t}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}}{1-\beta_{1}}\left\|{\bm{m}}_{t-1}\right\|\left\|\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}}\right)\right\|$ (191)
@@
$\leq\frac{2\left(\alpha_{t-1}-\alpha_{t}\right)^{2}\beta_{1}^{2}\left(1-\beta_{1}^{t-1}\right)G^{2}}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}}+\frac{2\alpha_{t}^{2}G^{2}}{\epsilon^{2}}+\frac{2\alpha_{t}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}\sqrt{1-\beta_{1}^{t-1}}G^{2}}{\left(1-\beta_{1}\right)\epsilon^{2}}.$ (193)